40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 27 18:03:47 2014, ericq, Update, Green Locking, uPDH Box Checkup uPDH_X.JPGuPDH_Y.JPGupdhTFs.pdfupdhNoises.pdfuPDH_Aug27.zip
    Reply  Thu Aug 28 17:34:20 2014, ericq, Update, Green Locking, uPDH Box Checkup TEK00006.PNG
       Reply  Fri Aug 29 17:28:07 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Fri Aug 29 23:53:35 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Tue Sep 2 20:02:06 2014, ericq, Update, Green Locking, uPDH Box Checkup 
       Reply  Wed Sep 3 00:17:22 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Wed Sep 3 20:38:40 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Fri Sep 5 04:07:44 2014, ericq, Update, Green Locking, X end uPDH Box Replaced 
                Reply  Fri Sep 5 16:13:00 2014, ericq, Update, Green Locking, new X Green PDH measurements 6x
                   Reply  Sat Sep 6 01:48:54 2014, rana, Update, Green Locking, new X Green PDH measurements 
                      Reply  Sat Sep 6 14:49:12 2014, ericq, Update, Green Locking, new X Green PDH measurements xGComparison.pdfxGspectra.pdf
                         Reply  Sun Sep 7 17:06:38 2014, rana, Update, Green Locking, new X Green PDH measurements 
                         Reply  Mon Sep 8 14:14:13 2014, Jenne, Update, Green Locking, new Y Green PDH measurements 6x
Message ID: 10460     Entry time: Fri Sep 5 16:13:00 2014     In reply to: 10457     Reply to this: 10463
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: new X Green PDH measurements 

 I measured the noise spectra and loop TF of the green PDH with the newly stuffed board. Unfortunately, I never took the noise below 100Hz of the previous box, so we can't see what has happened to the overall RMS, or more specifically, the RMS due to the pendulum resonance. All of these plots are in the boosted state, as that is how we intend to use the box. 

Here is the loop, which does not have quite as much margin as the y-arm, but 10dB of gain peaking is probably ok, since the RMS at 10s of kHz is not so important to ALS. (OL measured, CL inferred)  We see the 1/f shape from 1k to 50k or so, and 1/f^2 under 1k, as desired. 

xGbodes.pdf

Comparing in the in loop error signals, we see the effect from the increased gain from 100Hz to 10kHz. (Here is where I regret not looking at the low frequency spectrum two weeks ago)

pdhComparison.pdf

Finally, here is the noise breakdown. 

xGspectra.pdf

The error signal RMS is now dominated by the 1Hz peak. We have talked about using digital feedback for this, since we have the PDH error signal coming into an ADC, and can sum in a DAC signal into the servo output. This also lets us intelligently trigger a sub-10Hz boost once the PDH box locks itself. With a good boost, we maybe could bring the in-loop RMS of the error signal to under 1kHz. 


Something odd that Rana brought to my attention, however, is that my measurement and calibration indicates an RMS of ~5kHz, but the cavity pole should be something like 18kHz. If this is true, how can we be seeing stable power? This maybe means that my calibration is too many Hz per Volt.

I performed the calibration by creating a MIST model of the arm, and generating the PDH error signal on a demodulated PD, I then find the slope of Hz per arbitrary error signal unit. Then, looking at a scope trace, I match up the horn-to-horn voltage to the horn-to-horn arbitrary error signal units, which lets me finally find Hz per error signal volt. 

However, there is some qualitative difference in the shape between the simulated and observed error signals, namely, that the outer horns are larger than the inner horns in the real signal. 

sigSim.pdfscopeSweep.jpg

 

Does this matter? Is there something in my simulation that I can correct that would give a more accurate calibration? 

Data, plots, code, attached. 

Attachment 6: xG_Sep5.zip  653 kB  Uploaded Fri Sep 5 17:13:59 2014
ELOG V3.1.3-